# A Lie algebraic pattern behind logarithmic CFTs

### v2

@@ 1.4. Borel-Weil-Bott type theorem
$0\rightarrow W_{i,\lambda}(\mu)\rightarrow V_{\lambda}(\mu)\rightarrow W_{i,\sigma_{i}\ast\lambda}(\mu+\sigma_{i}\uparrow\lambda)\rightarrow 0$ (1.9)
@@
[Sug21, Lemma 4.10] For $i\in I$, $\sigma\in W$, $\lambda\in\Lambda$ such that $\ell(\sigma\sigma_{i})=\ell(\sigma)+1$ and $\sigma\ast\lambda\not\in\Lambda^{\sigma_{i}}$, we have short exact sequences of $P_{i}$-modules
@@
$0\rightarrow H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})\otimes W_{i,\sigma_{i}\sigma\ast\lambda}\rightarrow H^{1}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\rightarrow H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})\otimes W_{i,\sigma\ast\lambda}\rightarrow 0$
@@
Let us apply the long exact sequence $-H^{\bullet}(P_{i}\times_{B}\text{-})$ for (1.9) with
@@
[Sug21, Theorem 4.8] Let us assume that $V_{\lambda}$ is a $P_{i}$-module if $\lambda\in\Lambda^{\sigma_{i}}$. For $\lambda\in\Lambda$ and $\sigma\in W$ satisfying $\ell(\sigma_{i}\sigma)=\ell(\sigma)+1$ and $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$, we have an isomorphism
$H^{n}(G\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq H^{n+1}(G\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda)).$ (1.10)
@@
$H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\simeq\delta_{n,0}\mathbb{C}_{\sigma\uparrow\lambda},\quad H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)})=0$ (1.11)
@@
$H^{n}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq\delta_{n,0}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda).$ (1.12)
@@
$\sigma_{i}\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)),~{}\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)).$ (1.13)
@@
$H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})=0,~{}H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})\simeq\delta_{n,1}\mathbb{C}_{\sigma\uparrow\lambda},$ (1.14)
@@
$H^{n}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\simeq\delta_{n,1}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda).$ (1.15)
@@
$H^{a}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq\delta_{a,0}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda)\simeq H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))$ (1.16)
@@
$\begin{split}H^{a}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))&\simeq H^{a}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\otimes V_{\sigma\ast\lambda}\\ &\simeq H^{b}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})\otimes V_{\sigma_{i}\sigma\ast\lambda}\\ &\simeq H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\end{split}$ (1.17)
@@
$H^{1}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq H^{0}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\simeq 0.$ (1.18)
@@
1.5. Weyl type character formula
For a graded vector space $V=\bigoplus_{\Delta}V_{\Delta}$, $\dim_{\mathbb{C}}V_{\Delta}<\infty$, denote $\operatorname{ch}_{q}V$ the ($q$-)character
@@ 1.5. Weyl type character formula
(where $q$ and $\Delta$ can take multiple variables, i.e. $q=(q_{1},\dots,q_{n})$, $\Delta=(\Delta_{1},\dots,\Delta_{n})$ and $q^{\Delta}=q_{1}^{\Delta_{1}}\cdots q_{n}^{\Delta_{n}}$). In the same manner as [Sug21, Section 4.4], by combining the cohomology vanishing 666 Note that the Kempf’s vanishing theorem [Jan03, II,4] is proved under a weaker condition than Borel-Weil-Bott theorem [Jan03, II,5]. The same thing might hold in our case, but the verification of this is a future work (see also the footnote in Section 2.2). $H^{n>0}(G\times_{B}V_{\lambda})=0$ and the Atiyah-Bott localization formula [AB68], we obtain the Weyl-type character formula of the Feigin-Tipunin construction $H^{0}(G\times_{B}V_{\lambda})$. An application of the case $n=0$ in Theorem 1.1(2) will be discussed in Section 2.2.
@@
For $\lambda\in\Lambda$ satisfying a condition (strong), we have
@@
$\operatorname{ch}\mathcal{W}_{-\alpha+\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ(\alpha+\lambda^{\bullet})}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}$ (1.19)
@@ 2.1. Preliminary from VOSA
  • (translation coinvariance): $[T,a(z)]=\partial a(z)$,
@@
By abuse of notation, we simply write $V$ for the corresponding VSA. Other representation theoretic concepts (e.g., homomorphism, module,…) are defined in a common way and we omit it (for more detail, see [FB01]). For a VSA $V$, the parity auntomorphism $\iota_{V}\in\operatorname{Aut}(V)$ is defined by $\iota_{V}(a)=p(a)a$. Note that $\iota_{V_{1}\otimes V_{2}}=\iota_{V_{1}}\otimes\iota_{V_{2}}$.
@@
$[L_{m},L_{n}]=(m-n)L_{m+n}+\tfrac{m^{3}-m}{12}\delta_{m+n,0}c_{V},\ \ [L_{n},c_{V}]=0,\ \ (m,n\in\mathbb{Z}).$ (2.2)
For a VOSA-module $M$, we call $M$ is graded if there exists $a^{1},\dots,a^{n}\in V$ such that $L_{0},a^{1}_{(0)},\dots,a^{n}_{(0)}$ acts on $M$ semisimply and each eigenspace with respect to $(L_{0},a^{1}_{(0)},\dots,a^{n}_{(0)})$ is finite-dimensional, $\Delta_{a^{i}}=1$ and $L_{1}a^{i}=0$ for each $1\leq i\leq n$. We will use the notation
@@
$\operatorname{ch}_{q,z_{1},\dots,z_{n}}M=\operatorname{tr}_{M}q^{L_{0}-\tfrac{c}{24}}z_{1}^{a^{1}_{(0)}}\cdots z_{n}^{a^{n}_{(0)}},\ \ (\text{resp.\ $\operatorname{sch}_{q,z_{1},\dots,z_{n}}M=\sum_{x\in\{0,1\}}(-1)^{x}\operatorname{tr}_{M_{\bar{x}}}q^{L_{0}-\tfrac{c}{24}}z_{1}^{a^{1}_{(0)}}\cdots z_{n}^{a^{n}_{(0)}}$})$
A graded VOSA-module $M$ is CFT-type if there exists $m\in M_{\Delta,x_{1},\dots,x_{n}}$ for some $\Delta,x_{1},\dots,x_{n}$ such that $M_{\Delta,x_{1},\dots,x_{n}}=\mathbb{C}m$ and $M=U(V)m$. Note that a VOSA is always CFT-type because $a=a_{(-1)}\mathbf{1}$. If we do not consider $a^{1},\dots,a^{n}$, we will call the grading conformal grading. After Section 2.3, we only work with conformally-graded VOSAs and their modules.  [^7] On the other hand, for example in [CNS24], we consider the grading by $(L_{0},a^{1}_{(0)}:=h_{(0)})$ for $h\in\mathfrak{sl}_{2}\subset\hat{\mathfrak{sl}_{2}}$.
If we allow the exponent of $z$ in field expansion to be a rational number, then one can define the generalized V(O)SA in a similar manner (see [DL93] for more details). Let us introduce the following two operations that create another VOSA module from a given one.
@@
    (Contragredient dual) The contragredient dual $M^{\ast}$ of $M$ is defined as the restricted dual of $M$, i.e., $M^{\ast}=\bigoplus_{\Delta,x_{1},\dots,x_{n}}M_{\Delta,x_{1},\dots,x_{n}}^{*}$ with the following VO(S)A-module structure [FHL93]:
@@
    $\langle m^{\prime},L_{n}m\rangle=\langle L_{-n}m^{\prime},m\rangle,\ \ (a^{i}_{(0)})^{\dagger}=a^{i}_{(0)}.$ (2.3)
@@
    (Twisted module and spectral flow twist) For $g\in\operatorname{Aut}(V)$ with finite order $s\in\mathbb{Z}_{\geq 1}$. we can define $g$-twisted modules of $V$, which play an important role in orbifold conformal theory (see [Li96] for more details). Let us give a procedure called spectral flow twist which deform a given twisted $V$-module. Suppose that there exists $h\in V_{\overline{0}}$, called simple current element, and some $k\in\mathbb{C}$ such that
@@
    where $h_{(0)}$ acts semisimplely on $V$ and $\mathrm{Spec}(h_{(0)})$ refers to the set of eigenvalues of $h_{(0)}$. Then one can define Li’s Delta operator [Li97],
@@
    Let $M$ be a $g$-twisted $V$-module. Then $(M,Y_{M}(\Delta(z)\cdot,z))$ is a $gg_{h}$-twisted $V$-module, called spectral flow twist of $M$ and denote $\mathcal{S}_{h}(M)$, where $g_{h}=\exp(2\pi ih_{(0)})$ is an automorphism of $V$ with order $s$ (see [Li96, Proposition 5.4]). Note that if $M$ is irreducible, then so is $\mathcal{S}_{h}(M)$.
@@ 2.2. Simplicity theorem
This subsection will not be used until Section 3 and might be skipped once. In this subsection, we formalize the discussion in [Sug22] to enhance the perspective and flexibility of the discussion. 888 However, the authors expect that there should be a more concise way leading to the simplicity theorem in our case (ideally, they would like to include this subsection in Part I). One reason for this belief is that Theorem 1.1(2) has strong similarity to [Jan03, II,5], but in [Jan03, II,5] the simplicity theorem is proved as a direct corollary of the Borel-Weil-Bott theorem (i.e., no need to go through Kazhdan-Lusztig polynomials, etc.). It is a future work to study and develop the theory of shift system independently of VOA. Throughout this subsection, $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ is a shift system.
@@
[Sug22, Lemma 2.27, Corollary 3.3] For $\lambda,\lambda^{\prime}\in\Lambda$ satisfying the condition (strong), if
@@
as $G$-modules. For the naturality discussion, see [Sug22, Section 2]. ∎
@@
In [CNS24, Proposition 6.12], the isomorphisms in Lemma 2.4 hold up to the spectral flow twist $\mathcal{S}^{2}$. However, since $\mathcal{S}^{2}$ commutes with the $B$-action, it does not affect to the discussion and we omit it.
@@
    $W_{P}:=\bigoplus_{\begin{subarray}{c}\mu\in\Lambda\\ \mu_{\bullet}=0\end{subarray}}H^{0}(G\times_{B}V_{\mu})$ is a generalized CFT-type VOSA and $W_{P+\lambda_{\bullet}}:=\bigoplus_{\begin{subarray}{c}\mu\in\Lambda\\ \mu_{\bullet}=\lambda_{\bullet}\end{subarray}}H^{0}(G\times_{B}V_{\mu})$ is a graded $W_{P}$-module for each $\lambda\in\Lambda$. Furthermore, the natural $G$-action is in $\operatorname{Aut}(W_{P})$ and $W_{P}$ is an abelian intertwining algebra with abelian group $P/Q$ (see [McR20]).
If (2) holds, then so does (1) (but the converse is not true in general). If (1) (resp. (2)) holds, then $\mathcal{W}_{0}\simeq W_{Q}^{G}\simeq W_{P}^{G}$ is a vertex operator super subalgebra of $W_{Q}$ and $\mathcal{W}_{-\beta}$ is a $\mathcal{W}_{0}$-module for any $\beta\in P_{+}\cap Q$ (resp. for any $\beta\in P_{+}$) and $\lambda\in\Lambda$. In the same manner as [Sug22, Section 3.2], we obtain the following.
@@
We will omit the detail of the proof (see [Sug22, Section 3.2]). Let us consider the case (2). By Lemma 2.4, $W_{P}$ is self-dual. Then for the non-degenerate $W_{P}$-invariant bilinear form $\langle\ ,\ \rangle\colon W_{P}\times W_{P}\rightarrow\mathbb{C}$, we have
@@
where $a\in W_{P}$ and $a^{\ast}\in W_{P}\simeq W_{P}^{\ast}$ by abuse of notation. Since $W_{P}$ is CFT-type, we have $(a^{\ast}_{(-1)})^{\dagger}a=\mathbf{1}$ and thus $W_{P}$ is simple. By [McR20, Theorem 3.2], $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module for any $\beta\in P_{+}$. Finally, by [McR20, Proposition 2.26], each $W_{Q+\lambda^{\bullet}}$ is simple as $W_{Q}$-module. ∎
In the same manner as [Sug22, Lemma 3.5, 3.22] or [CNS24, Theorem 6.15], we have the following.

### v3

@@ 1.4. Borel–Weil–Bott-type theorem
(1.10) $0\rightarrow W_{i,\lambda}(\mu)\rightarrow V_{\lambda}(\mu)\rightarrow W_{i,\sigma_{i}\ast\lambda}(\mu+\sigma_{i}\uparrow\lambda)\rightarrow 0.$
@@
[45, Lemma 4.10] For $i\in I$, $\sigma\in W$, $\lambda\in\Lambda$ such that $\ell(\sigma\sigma_{i})=\ell(\sigma)+1$ and $\sigma\ast\lambda\not\in\Lambda^{\sigma_{i}}$, we have short exact sequences of $P_{i}$-modules
@@
$0$ $\rightarrow H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})\otimes W_{i,\sigma_{i}\sigma\ast\lambda}$
$\rightarrow H^{1}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))$
$\rightarrow H^{1}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})\otimes W_{i,\sigma\ast\lambda}\rightarrow 0.$
@@
Let us apply the long exact sequence $H^{\bullet}(P_{i}\times_{B}\text{-})$ for (1.10) with
@@
[45, Theorem 4.8] Let us assume that $V_{\lambda}$ is a $P_{i}$-module if $\lambda\in\Lambda^{\sigma_{i}}$. For $\lambda\in\Lambda$ and $\sigma\in W$ satisfying $\ell(\sigma_{i}\sigma)=\ell(\sigma)+1$ and $(\sigma\uparrow\lambda,\alpha_{i}^{\vee})=0$, we have an isomorphism
(1.11) $H^{n}(G\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq H^{n+1}(G\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda)).$
@@
(1.12) $H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\simeq\delta_{n,0}\mathbb{C}_{\sigma\uparrow\lambda},\quad H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)})=0$
@@
(1.13) $H^{n}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq\delta_{n,0}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda).$
@@
(1.14) $\sigma_{i}\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma\ast\lambda)),~\sigma\uparrow\lambda=\sigma_{i}\circ(\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)).$
@@
(1.15) $H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})=0,~H^{n}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda+\sigma_{i}\uparrow(\sigma_{i}\sigma\ast\lambda)})\simeq\delta_{n,1}\mathbb{C}_{\sigma\uparrow\lambda},$
@@
(1.16) $H^{n}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\simeq\delta_{n,1}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda).$
@@
(1.17) $H^{a}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq\delta_{a,0}W_{i,\sigma\ast\lambda}(\sigma\uparrow\lambda)\simeq H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))$
@@
(1.18) $\begin{split}H^{a}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))&\simeq H^{a}(P_{i}\times_{B}\mathbb{C}_{\sigma\uparrow\lambda})\otimes V_{\sigma\ast\lambda}\\ &\simeq H^{b}(P_{i}\times_{B}\mathbb{C}_{\sigma_{i}\sigma\uparrow\lambda})\otimes V_{\sigma_{i}\sigma\ast\lambda}\\ &\simeq H^{b}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\end{split}$
@@
(1.19) $H^{1}(P_{i}\times_{B}V_{\sigma\ast\lambda}(\sigma\uparrow\lambda))\simeq H^{0}(P_{i}\times_{B}V_{\sigma_{i}\sigma\ast\lambda}(\sigma_{i}\sigma\uparrow\lambda))\simeq 0.$
@@
1.5. Weyl-type character formula
For a graded vector space $V=\bigoplus_{\Delta}V_{\Delta}$, $\dim_{\mathbb{C}}V_{\Delta}<\infty$, denote
@@ 1.5. Weyl-type character formula
the character (where $q$ and $\Delta$ can take multiple variables, i.e. $q=(q_{i})_{1\leq i\leq n}$, $\Delta=(\Delta_{i})_{1\leq i\leq n}$, $q^{\Delta}=\Pi_{i=1}^{n}q^{\Delta_{i}}$). In the same manner as [45, Section 4.4], by combining the cohomology vanishing777 Note that the Kempf’s vanishing theorem [32, II,4] is proved under a weaker condition than BWB theorem [32, II,5]. The same situation might hold in our case, but the verification of this is a future work (see also the footnote in Section 2.2). $H^{n>0}(G\times_{B}V_{\lambda})=0$ and the Atiyah–Bott localization formula [3], we obtain the Weyl-type character formula of the FT construction $H^{0}(G\times_{B}V_{\lambda})$. An application of the case $n=0$ in Theorem 1.1(2) will be discussed in Section 2.2.
@@
For $\lambda\in\Lambda$ satisfying the condition (strong), we have
@@
(1.20) $\operatorname{ch}\mathcal{W}_{-\alpha+\lambda}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\lambda}^{h=\sigma\circ(\alpha+\lambda^{\bullet})}=\sum_{\sigma\in W}(-1)^{l(\sigma)}\operatorname{ch}_{q}V_{\sigma\ast\lambda}^{h=\alpha+\lambda^{\bullet}-\sigma\uparrow\lambda}$
@@ 2.1. Preliminary from VOSA
  • (translation coinvariance): $T,a(z)]=\partial a(z)$,
@@
By abuse of notation, we simply write $V$ for the corresponding VSA. Other representation theoretic concepts (e.g., homomorphism, module,…) are defined in a common way and we omit it (for more detail, see [20]). For a VSA $V$, the parity automorphism $\iota_{V}\in\operatorname{Aut}(V)$ is defined by $\iota_{V}(a)=p(a)a$. Note that $\iota_{V_{1}\otimes V_{2}}=\iota_{V_{1}}\otimes\iota_{V_{2}}$.
@@
(2.2) $[L_{m},L_{n}]=(m-n)L_{m+n}+\tfrac{m^{3}-m}{12}\delta_{m+n,0}c_{V},\ \ [L_{n},c_{V}]=0,\ \ (m,n\in\mathbb{Z}).$
[…]
@@
[…]
A graded VOSA-module $M$ is CFT-type if there exists $m\in M_{\Delta,x_{1},\dots,x_{n}}$ for some $\Delta,x_{1},\dots,x_{n}$ such that $M_{\Delta,x_{1},\dots,x_{n}}=\mathbb{C}m$ and $M=U(V)m$. Note that a VOSA is always CFT-type because $a=a_{(-1)}\mathbf{1}$. If we do not consider $a^{1},\dots,a^{n}$, we will call the grading conformal grading. After Section 2.3, we only work with conformally-graded VOSAs and their modules. [^8] On the other hand, for example in [14], we consider the grading by $(L_{0},a^{1}_{(0)}:=h_{(0)})$ for $h\in\mathfrak{sl}_{2}\subset\hat{\mathfrak{sl}_{2}}$.
If we allow the exponent of $z$ in field expansion to be a rational number, then one can define the generalized V(O)SA in a similar manner (see [18] for more details). Let us introduce the following two operations that create another VOSA module from a given one.
@@
    (Contragredient dual) The contragredient dual $M^{\ast}$ of $M$ is defined as the restricted dual of $M$, i.e., $M^{\ast}=\bigoplus_{\Delta,x_{1},\dots,x_{n}}M_{\Delta,x_{1},\dots,x_{n}}^{*}$ with the following VO(S)A-module structure [25]:
@@
    (2.3) $\langle m^{\prime},L_{n}m\rangle=\langle L_{-n}m^{\prime},m\rangle,\ \ (a^{i}_{(0)})^{\dagger}=a^{i}_{(0)}.$
@@
    (Twisted module and spectral flow twist) For $g\in\operatorname{Aut}(V)$ with finite order $s\in\mathbb{Z}_{\geq 1}$, we can define $g$-twisted modules of $V$, which play an important role in orbifold conformal theory (see [39] for more details). Let us give a procedure called spectral flow twist which deform a given twisted $V$-module. Suppose that there exists $h\in V_{\overline{0}}$, called simple current element, and some $k\in\mathbb{C}$ such that
@@
    where $h_{(0)}$ acts semisimplely on $V$ and $\mathrm{Spec}(h_{(0)})$ refers to the set of eigenvalues of $h_{(0)}$. Then one can define Li’s Delta operator [40],
@@
    Let $M$ be a $g$-twisted $V$-module. Then $(M,Y_{M}(\Delta(z)\cdot,z))$ is a $gg_{h}$-twisted $V$-module, called spectral flow twist of $M$ and denote $\mathcal{S}_{h}(M)$, where $g_{h}=\exp(2\pi ih_{(0)})$ is an automorphism of $V$ with order $s$ (see [39, Proposition 5.4]). Note that if $M$ is irreducible, then so is $\mathcal{S}_{h}(M)$.
@@ 2.2. Simplicity theorem
This subsection will not be used until Section 3 and might be skipped once. In this subsection, we formalize the discussion in [46] to enhance the perspective and flexibility of the discussion.999 However, the authors expect that there should be a more concise way leading to the simplicity theorem in our case (ideally, they would like to include this subsection in Part I). One reason for this belief is that Theorem 1.1(2) has strong similarity to [32, II,5], but in [32, II,5] the simplicity theorem is proved as a direct corollary of the BWB theorem (i.e., no need to go through Kazhdan–Lusztig polynomials, etc.). It is a future work to study and develop the theory of shift system independently of VOA. Throughout this subsection, $(\Lambda,\uparrow,\{V_{\lambda}\}_{\lambda\in\Lambda})$ is a shift system.
@@
[46, Lemma 2.27, Corollary 3.3] For $\lambda,\lambda^{\prime}\in\Lambda$ satisfying the condition (strong), if
@@
as $G$-modules. For the naturality discussion, see [46, Section 2]. ∎
@@
In [14, Proposition 6.12], the isomorphisms in Lemma 2.4 hold up to the spectral flow twist $\mathcal{S}^{2}$. However, since $\mathcal{S}^{2}$ commutes with the $B$-action, it does not affect to the discussion and we omit it.
@@
    $W_{P}:=\bigoplus_{\begin{subarray}{c}\mu\in\Lambda\\ \mu_{\bullet}=0\end{subarray}}H^{0}(G\times_{B}V_{\mu})$ is a generalized CFT-type VOSA and $W_{P+\lambda_{\bullet}}:=\bigoplus_{\begin{subarray}{c}\mu\in\Lambda\\ \mu_{\bullet}=\lambda_{\bullet}\end{subarray}}H^{0}(G\times_{B}V_{\mu})$ is a graded $W_{P}$-module for each $\lambda\in\Lambda$. Furthermore, the natural $G$-action is in $\operatorname{Aut}(W_{P})$ and $W_{P}$ is an abelian intertwining algebra with abelian group $P/Q$ (see [43]).
If (2) holds, then so does (1) (but the converse does not in general). If (1) (resp. (2)) holds, then $\mathcal{W}_{0}\simeq W_{Q}^{G}\simeq W_{P}^{G}$ is a vertex operator super subalgebra of $W_{Q}$ and $\mathcal{W}_{-\beta}$ is a $\mathcal{W}_{0}$-module for any $\beta\in P_{+}\cap Q$ (resp. for any $\beta\in P_{+}$) and $\lambda\in\Lambda$. In the same manner as [46, Section 3.2], we obtain the following.
@@
We will omit the detail of the proof (see [46, Section 3.2]). Let us consider the case (2). By Lemma 2.4, $W_{P}$ is self-dual. Then for the non-degenerate $W_{P}$-invariant bilinear form $\langle\ ,\ \rangle\colon W_{P}\times W_{P}\rightarrow\mathbb{C}$, we have
@@
where $a\in W_{P}$ and $a^{\ast}\in W_{P}\simeq W_{P}^{\ast}$ by abuse of notation. Since $W_{P}$ is CFT-type, we have $(a^{\ast}_{(-1)})^{\dagger}a=\mathbf{1}$ and thus $W_{P}$ is simple. By [43, Theorem 3.2], $\mathcal{W}_{-\beta}$ is simple as $\mathcal{W}_{0}$-module for any $\beta\in P_{+}$. Finally, by [43, Proposition 2.26], each $W_{Q+\lambda^{\bullet}}$ is simple as $W_{Q}$-module. ∎
In the same manner as [46, Lemma 3.5, 3.22] or [14, Theorem 6.15], we have the following.